A propeller is modeled as five identical uniform rods extending radially from its axis. The length and mass of each rod are 0.777 m and 2.67 kg, respectively. When the propellor rotates at 573 rpm (revolutions per minute), what is its rotational kinetic energy?
Answer #1

A single rod's  moment of inertia rotating about its end  is I=(1/3)mR2   

Kinetic energy is Ke= 0.5 Iω2  

I= ΣI = 5(1/3)mR2   

ω=573x2π/60= 60.0 rad/s

Now 

Ke= 0.5x (5/3) mR2   ω2  =

Ke= (5/6) 2.67 x (0.777)2  (60.0)2 

Ke= 4836 J  (or 4840 J)

Please let me know if you have any questions